Aqueous hydrochloric acid HCl will react with solid sodium hydroxide NaOH to produce aqueous sodium chloride NaCl and liquid water H2O. Suppose 27. g of hydrochloric acid is mixed with 10.1 g of sodium hydroxide. Calculate the maximum mass of water that could be produced by the chemical reaction. Round your answer to 3 significant digits.
1
Expert's answer
2019-10-17T07:30:10-0400

given reaction is-


"HCl(aq)+NaOH(aq)\to NaCl(aq)+H_2O"

moles of HCl taken =

"=\frac{27}{36.5}=0.7397\ mole"

"\frac{}{}" moles of NaOH taken

"=\frac{10.1}{40}=0.2525\ mole"

So NaOH is the limiting reagent here

then

moles of water formed = moles of NaOH taken


"=0.2525\ mole\ H_2O=0.2525\times 18=4.54\ gm\ H_2O"